What is the molarity of HCl(aq) if 17.65 mL of HCl completely reacts with 10.00 mL of 0.250 M K2CO3
1
Expert's answer
2019-09-23T03:46:47-0400

The reaction proceeds according to the equation:

"2HCl + K_2CO_3 = 2KCl + CO_2 + H_2O;"

The total amount of moles of potassium carbonate is:

"n(K_2CO_3 ) = 0.250 M * 0.01 L = 0.0025 mol;"

The amount of moles of HCl is twice bigger than amount of potassium carbonate moles (quotient 2 near the HCl in equation), so there were 0.0025*2=0.005 moles of HCl in reaction.

Now we can calculate the HCl solution molarity:

"C_M(HCl) = \frac {n(HCl)} {V(HCl)} = \frac {0.005 moles} {0.01765 L} = 0.283M"

Answer: the molarity of the HCl is 0.283M.